Use the sample data and confidence level given below to complete parts​ (a) through​ (d). A research institute poll asked respondents if they felt vulnerable to identity theft. In the​ poll, n equals 939 and x equals 505 who said​ "yes." Use a 99 % confidence level. ​a) Find the best point estimate of the population proportion p. nothing ​(Round to three decimal places as​ needed.) ​b) Identify the value of the margin of error E. Eequals nothing ​(Round to three decimal places as​ needed.) ​c) Construct the confidence interval. nothingless than p less than nothing ​(Round to three decimal places as​ needed.) ​d) Write a statement that correctly interprets the confidence interval. Choose the correct answer below. A. 99​% of sample proportions will fall between the lower bound and the upper bound. B. One has 99​% confidence that the sample proportion is equal to the population proportion. C. One has 99​% confidence that the interval from the lower bound to the upper bound actually does contain the true value of the population proportion. D. There is a 99​% chance that the true value of the population proportion will fall between the lower bound and the upper bound.
Answer #1

A)

Point estimate = 505 / 939 = 0.538

B)

Margin of error = Z/2 * sqrt [ ( 1 - ) / n ]

= 2.576 * sqrt [ 0.538 * ( 1 - 0.538) / 939 ]

= 0.042

c)

99% confidence interval is

- E < p < + E

0.538 - 0.042 < p < 0.538 + 0.042

0.496 < p < 0.580

d)

One has 99​% confidence that the interval from the lower bound to the upper bound actually does contain

the true value of the population proportion.